a proton of mass m and charge q enters a magnetic field B with velocity v at an angle x with the direction of B what will be the radius of curvature of resulting path

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Force acting on proton=\left| q(\vec{v}\times \vec{B})\right|=qvBsinx
This force causes the centripetal acceleration of the proton. Hence

qvBsinx=\dfrac{m(v sinx )^2}{R}
\implies R=\dfrac{mv sinx}{qB}